# Shift-cyclicity in Analytic Function Spaces

### v1

@@ 3.1.1. The Dirichlet space $\mathcal{D}$
    $\mathcal{D}\subsetneq H^{2}(\mathbb{D})$ [^7] In fact, $\mathcal{D}\subsetneq\bigcap_{p<\infty}H^{p}(\mathbb{D})$, $H^{\infty}(\mathbb{D})\not\subset\mathcal{D}$ and $\mathcal{D}\not\subset H^{\infty}(\mathbb{D})$ [[, see]Exercise 1.1 (7)]EKMR14.. Thus, functions in $\mathcal{D}$ have radial limits a.e. on $\mathbb{T}$ and exhibit an inner-outer factorization.
@@
    If $f^{\prime}\in H^{2}(\mathbb{D})$ then $f\in\mathcal{D}$. The function
    $f(z):=\sum_{n\in\mathbb{N}}\frac{z^{n}}{n}$
    shows that the converse is not true.
Things get complicated from here on. For starters, one can check that if $\{w_{n}:n\in\mathbb{N}\}$ is the zero set of a Blaschke product $B$, then $f(z)=(1-z)^{2}B(z)$ is such that $f^{\prime}\in H^{\infty}(\mathbb{D})$. Thus, $(4)$ above shows that $\{w_{n}:n\in\mathbb{N}\}$ is the zero set of some $f\in\mathcal{D}$. Interestingly, infinite Blaschke products do not lie in $\mathcal{D}$. In fact, the only inner functions that lie in $\mathcal{D}$ are finite Blaschke products (this follows from a result of Carleson [Car60]; see also [EKMR14, Theorem 4.1.3]). Determining all the zero sets of functions in $\mathcal{D}$ is still an open problem.
Next, we note that $H^{\infty}(\mathbb{D})$ is not the multiplier algebra of $\mathcal{D}$. In fact, the multiplier algebra $\mathcal{M}(\mathcal{D})$ does not seem to have a simple description in terms of integral or functional properties of its members. Since $\mathcal{D}$ does not contain every inner function, we know that $H^{\infty}(\mathbb{D})\not\subset\mathcal{D}$ and it is not difficult to check that
@@
but [EKMR14, Theorem 5.1.6] shows that
@@
Lastly, every shift-invariant subspace of $\mathcal{D}$ is of the form $S[\varphi]$ for some $\varphi\in\mathcal{M}(\mathcal{D})$ [^8] A corollary of this fact is that the zero sets of members in $\mathcal{D}$ and $\mathcal{M}(\mathcal{D})$ are the same. [[, see]Theorem 2 (c)]RS88, but we do not have a good description of all such $\varphi$’s. This does tell us that MC generalizes to $\mathcal{D}$ with $\mathcal{M}(\mathcal{D})$ instead of $H^{\infty}(\mathbb{D})$. [EKMR14, Theorem 9.1.3] shows that if $f\in\mathcal{D}$, then
@@
Consequently, if $f$ is shift-cyclic in $\mathcal{D}$, then it must be outer. However, just like the Hardy spaces for $d>1$, there are outer functions that are not shift-cyclic [[, see]Theorem 9.2.8]EKMR14. It turns out that in order to understand shift-cyclicity in $\mathcal{D}$ we have one more necessary condition to consider alongside outerness.
@@ 3.1.2. Logarithmic capacity and shift-cyclicity
A set $\Gamma\subset\mathbb{T}$ is said to be capacitable if $\operatorname{cap}_{*}(\Gamma)=\operatorname{cap}^{*}(\Gamma)$, in which case we simply write $\operatorname{cap}(\Gamma)$ for either. A classical result of Choquet [Cho54] shows that every Borel subset of $\mathbb{T}$ is capacitable. The capacity of a Borel set $\Gamma\subset\mathbb{T}$ is intimately connected with its (normalized) Lebesgue measure $|\Gamma|$. For instance, [EKMR14, Theorem 2.4.5] shows that if $|\Gamma|>0$ then
@@
It follows that if $\operatorname{cap}(\Gamma)=0$ then $|\Gamma|=0$ as well. However, the Cantor middle thirds set $C_{1/3}$ is such that $|C_{1/3}|=0$ but $\operatorname{cap}(C_{1/3})>0$ [[, see]Theorem 2.3.5]EKMR14. For this reason, the following result of Beurling [Beu40] highlights an important deviation of $\mathcal{D}$ from the Hardy spaces. Recall that a property holds quasi everywhere (q.e.) if it holds everywhere except on a set $\Gamma$ with $\operatorname{cap}^{*}(\Gamma)=0$.
@@
The following result of Brown and Shileds [BS84, Theorem 5] shows how logarithmic capacity relates to shift-cyclicity.
@@
As a result of this theorem, we arrive at the famous Brown–Shields conjecture (appeared as Question 12 in [BS84]), which is still not resolved.
@@
A compilation of sufficient conditions for the shift-cyclicity of a given $f\in\mathcal{D}$ can be found in [EKMR14, Chapter 9], but, in general, the question is still open. We end this section with some notable results about shift-cyclicity in $\mathcal{D}$.
@@
    A polynomial is shift-cyclic in $\mathcal{D}$ if and only if it is non-vanishing on $\mathbb{D}$ [[, see]Theorem 9.2.1]EKMR14.
@@
    If Problem 3.3 can be resolved for $f\in\mathcal{M}(\mathcal{D})$, then it can be resolved entirely [[, see]Corollary 2.5.2]Sam22.
@@ 3.1.3. Dirichlet-type spaces on $\mathbb{D}$
is the normalized area measure on $\mathbb{D}$ [[, see]Lemma 2]Tay66. Note that for $t=0$ and $t=1$, we recover $H^{2}(\mathbb{D})$ and $\mathcal{D}$ respectively. Another important space in this family is the Bergman space $\mathcal{D}_{-1}=L^{2}_{a}(\mathbb{D})$ which will turn up frequently in Section 4. Also, note that
@@
For $t>1$, functions in $\mathcal{D}_{t}$ are continuous up to $\overline{\mathbb{D}}$. In fact, $\mathcal{D}_{t}$ is an algebra under point-wise multiplication and, as a result, $\mathcal{D}_{t}$ is its own multiplier algebra. It is then not so difficult to check that $f$ is shift-cyclic in $\mathcal{D}_{t}$ if and only if it is non-vanishing on $\overline{\mathbb{D}}$ [[, see]pp. 274]BS84. For $t\leq 0$, the above equivalent integral norm shows that the multiplier algebra of $\mathcal{D}_{t}$ is $H^{\infty}(\mathbb{D})$, whereas the situation is much more complicated for $0<t\leq 1$. Also, for $t<0$, there are singular inner functions that turn out to be shift-cyclic in $\mathcal{D}_{t}$. Korenblum [Kor81] classified all shift-cyclic inner functions in $L^{2}_{a}(\mathbb{D})$. With these properties in mind, we highlight the following results from [BS84] (also see [BCLSS151] for a modern treatment of Dirichlet-type spaces on $\mathbb{D}$).
@@
    If $f\in H^{2}(\mathbb{D})$ [^9] This result is valid for all $f\in H^{1}(\mathbb{D})$ as well, since one can check that $H^{1}(\mathbb{D})\subset L^{2}_{a}(\mathbb{D})$. then it is shift-cyclic in $L^{2}_{a}(\mathbb{D})$ if and only if its inner part is shift-cyclic in $L^{2}_{a}(\mathbb{D})$ (which are determined in [Kor81]).
@@
Lastly, we remark that a polynomial is shift-cyclic in $\mathcal{D}_{t}$ with $t\leq 1$ if and only if it is non-vanishing on $\mathbb{D}$. This follows from the first part of (4) above, combined with [EKMR14, Theorem 9.2.1]. As noted before, a polynomial is shift-cyclic in $\mathcal{D}_{t}$ with $t>1$ if and only if it is non-vanishing on $\overline{\mathbb{D}}$. We saw with the Hardy spaces over $\mathbb{D}^{d}$ that determining the shift-cyclic polynomials is tractable. We use this as motivation to introduce a generalization of $\mathcal{D}_{t}$ over both $\mathbb{D}^{d}$ and $\mathbb{B}_{d}$, and remark that shift-cyclic polynomials have been classified for $d=2$.
@@ 3.2. On the unit polydisk $\mathbb{D}^{d}$
Moreover, we recover the Bergman, Hardy and Dirichlet spaces on $\mathbb{D}^{d}$ by taking $t=-1,0$ and $1$ respectively. For $t>1$, it can also be checked that $\mathcal{D}_{t}(\mathbb{D}^{d})$ is a subalgebra of functions in $\operatorname{Hol}(\mathbb{D}^{d})$ that are continuous up to $\overline{\mathbb{D}^{d}}$, and therefore, is its own multiplier algebra. Also, just like in the one variable case, a function $f$ is shift-cyclic in $\mathcal{D}_{t}(\mathbb{D}^{d})$ with $t>1$ if and only if it is non-vanishing on $\overline{\mathbb{D}^{d}}$ [[, see]Section 1.1]BKKLSS16.
While we can extend some of the general properties mentioned in the previous section to all $d>1$, one only needs to look at the polynomials to realize that classifying all shift-cyclic functions is further beyond the scope for these spaces than one might anticipate. In fact, we only have a complete classification of shift-cyclic polynomials for $d=2$. Let $\mathcal{Z}_{\Gamma}(P)$ denote the set of zeros of a polynomial $P\in\mathcal{P}_{d}$ on some $\Gamma\subseteq\mathbb{C}^{d}$. The following is a joint result of Bénéteau, Condori, Knese, Kosiński, Liaw, Seco and Sola [BCLSS152, BKKLSS16] [^10] See also [KKRS19] for a similar result for the anisotropic Dirichlet-type spaces on $\mathbb{D}^{2}$..
@@
If $P\in\mathcal{P}_{d}$ with $\mathcal{Z}_{\mathbb{D}^{d}}(P)=\emptyset$, then it is shift-cyclic in $\mathcal{D}_{t}(\mathbb{D}^{2})$ for
@@
    $t>1$ if and only if $\mathcal{Z}_{\mathbb{T}^{d}}(P)=\emptyset$.
There are a few things to unpack here. First, the proof of this theorem relies on something called a determinantal representation for polynomials that are non-vanishing on $\mathbb{D}^{2}$, which may not always exist for $d>2$ [Kne101, Kne102]. Next, this result clearly indicates how boundary zeros are important to study shift-cyclicity of polynomials. In particular, the authors use a notion of capacity called the Riesz $t$-capacity $\operatorname{cap}_{t}$ for the space $\mathcal{D}_{t}(\mathbb{D}^{2})$ to show that if $\operatorname{cap}_{t}(\mathcal{Z}_{\mathbb{T}^{2}}(P))$ is sufficiently large then $P$ cannot be shift-cyclic, which settles the case $t\leq 1/2$.
Arguably, the most interesting aspect of Theorem 3.4 is the case $1/2<t\leq 1$, where finiteness of the boundary zero set is crucial to determining whether a polynomial is shift-cyclic or not. The authors achieve this via the use of tools from semi-analytic geometry. More specifically, they use an inequality of Łojasiewicz [Loj59], which states that if $f$ is a non-zero real analytic function on some open set $U\subset\mathbb{R}^{N}$ with $N\in\mathbb{N}$ whose zero set $\mathcal{Z}_{U}(f)$ in $U$ is non-empty then, for every compact $V\subset U$, there are constants $C_{V}>0$ and $n_{V}\in\mathbb{N}$ such that
@@
Applying this inequality to $|P|^{2}$, where $P$ is a non-vanishing polynomial on $\mathbb{D}^{2}$ with finitely many zeros in $\mathbb{T}^{2}$, allows us to compare $P$ with other polynomials that are shift-cyclic [[, see]Lemma 3.3]BKKLSS16. This further enables us to determine the shift-cyclicity of such $P$ in $\mathcal{D}_{t}(\mathbb{D}^{2})$ when $t\leq 1$.
In the following section, we note that the Łojasiewicz inequality helps generalize this result to a similar family of spaces over $\mathbb{B}_{2}$, and also allows us to go beyond $d=2$ in this case.

### v3

@@ 3.1.1. The Dirichlet space $\mathcal{D}$
    $\mathcal{D}\subsetneq H^{2}(\mathbb{D})$ [^7] In fact, $\mathcal{D}\subsetneq\bigcap_{p<\infty}H^{p}(\mathbb{D})$, $H^{\infty}(\mathbb{D})\not\subset\mathcal{D}$ and $\mathcal{D}\not\subset H^{\infty}(\mathbb{D})$ (see [35, Exercise 1.1 (7)]).. Thus, functions in $\mathcal{D}$ have radial limits a.e. on $\mathbb{T}$ and exhibit an inner-outer factorization.
@@
    An inequality of Hardy (see [33, the corollary to Theorem 3.15]) shows that if $f^{\prime}\in H^{1}(\mathbb{D})$ then $f\in\mathcal{D}$. The function $f(z)=\sum_{n}\frac{z^{n}}{n}$ shows that $f$ need not lie in $\mathcal{D}$ even if $f^{\prime}\in H^{p}(\mathbb{D})$ for all $p<1$.
Things get complicated from here on. For starters, one can check that if $\{w_{n}:n\in\mathbb{N}\}$ is the zero set of a Blaschke product $B$, then $f(z)=(1-z)^{2}B(z)$ is such that $f^{\prime}\in H^{\infty}(\mathbb{D})$. Thus, $(4)$ above shows that $\{w_{n}:n\in\mathbb{N}\}$ is the zero set of some $f\in\mathcal{D}$. Interestingly, infinite Blaschke products do not lie in $\mathcal{D}$. In fact, the only inner functions that lie in $\mathcal{D}$ are finite Blaschke products (this follows from a result of Carleson [25]; see also [35, Theorem 4.1.3]). Determining all the zero sets of functions in $\mathcal{D}$ is still an open problem.
Next, we note that $H^{\infty}(\mathbb{D})$ is not the multiplier algebra of $\mathcal{D}$888$\mathcal{M}(\mathcal{D})$ is not even algebraically isomorphic to $H^{\infty}(\mathbb{D})$ (see [31, Corollary 7.4]).. In fact, the multiplier algebra $\mathcal{M}(\mathcal{D})$ does not seem to have a simple description in terms of integral or functional properties of its members. Since $\mathcal{D}$ does not contain every inner function, we know that $H^{\infty}(\mathbb{D})\not\subset\mathcal{D}$ and it is not difficult to check that
@@
but [35, Theorem 5.1.6] shows that
@@
Lastly, every shift-invariant subspace of $\mathcal{D}$ is of the form $S[\varphi]$ for some $\varphi\in\mathcal{M}(\mathcal{D})$ [^9] A corollary of this fact is that the zero sets of members in $\mathcal{D}$ and $\mathcal{M}(\mathcal{D})$ are the same. (see [89, Theorem 2 (c)]), but we do not have a good description of all such $\varphi$’s. This does tell us, though, that MC generalizes to $\mathcal{D}$ with $\mathcal{M}(\mathcal{D})$ instead of $H^{\infty}(\mathbb{D})$. [35, Theorem 9.1.3] shows that if $f\in\mathcal{D}$, then
@@
Consequently, if $f$ is shift-cyclic in $\mathcal{D}$, then it must be outer. However, just like the Hardy spaces for $d>1$, there are outer functions that are not shift-cyclic (see [35, Theorem 9.2.8]). It turns out that in order to understand shift-cyclicity in $\mathcal{D}$ we have one more necessary condition to consider alongside outerness.
@@ 3.1.2. Logarithmic capacity and shift-cyclicity
A set $\Gamma\subset\mathbb{T}$ is said to be capacitable if $\operatorname{cap}_{*}(\Gamma)=\operatorname{cap}^{*}(\Gamma)$, in which case we simply write $\operatorname{cap}(\Gamma)$ for either. A classical result of Choquet [27] shows that every Borel subset of $\mathbb{T}$ is capacitable. The capacity of a Borel set $\Gamma\subset\mathbb{T}$ is intimately connected to its (normalized) Lebesgue measure $|\Gamma|$. For instance, [35, Theorem 2.4.5] shows that if $|\Gamma|>0$ then
@@
It follows that if $\operatorname{cap}(\Gamma)=0$ then $|\Gamma|=0$ as well. However, the Cantor middle thirds set $C_{1/3}$ is such that $|C_{1/3}|=0$ but $\operatorname{cap}(C_{1/3})>0$ (see [35, Theorem 2.3.5]). For this reason, the following result of Beurling [18] highlights an important deviation of $\mathcal{D}$ from the Hardy spaces. Recall that a property holds quasi everywhere (q.e.) if it holds everywhere except on a set $\Gamma$ with $\operatorname{cap}^{*}(\Gamma)=0$.
@@
The following result of Brown and Shileds [24, Theorem 5] shows how logarithmic capacity relates to shift-cyclicity.
@@
As a result of this theorem, we arrive at the famous Brown–Shields conjecture (appeared as Question 12 in [24]), which is still not resolved.
@@
A compilation of sufficient conditions for the shift-cyclicity of a given $f\in\mathcal{D}$ can be found in [35, Chapter 9], but, in general, the question is still open. We end this section with some notable results about shift-cyclicity in $\mathcal{D}$.
@@
    A polynomial is shift-cyclic in $\mathcal{D}$ if and only if it is non-vanishing on $\mathbb{D}$ (see [35, Theorem 9.2.1]).
@@
    If Problem 3.3 can be resolved for $f\in\mathcal{M}(\mathcal{D})$, then it can be resolved entirely (see [97, Corollary 2.5.2]).
@@ 3.1.3. Dirichlet-type spaces on $\mathbb{D}$
is the normalized area measure on $\mathbb{D}$ (see [103, Lemma 2]). Note that for $t=0$ and $t=1$, we recover $H^{2}(\mathbb{D})$ and $\mathcal{D}$ respectively. Another important space in this family is the Bergman space $\mathcal{D}_{-1}=L^{2}_{a}(\mathbb{D})$ which will turn up frequently in Section 4. Also, note that
@@
For $t>1$, functions in $\mathcal{D}_{t}$ are continuous up to $\overline{\mathbb{D}}$. In fact, $\mathcal{D}_{t}$ is an algebra under point-wise multiplication and, as a result, $\mathcal{D}_{t}$ is its own multiplier algebra. It is then not so difficult to check that $f$ is shift-cyclic in $\mathcal{D}_{t}$ if and only if it is non-vanishing on $\overline{\mathbb{D}}$ (see [24, pp. 274]). For $t\leq 0$, the above equivalent integral norm shows that the multiplier algebra of $\mathcal{D}_{t}$ is $H^{\infty}(\mathbb{D})$, whereas the situation is much more complicated for $0<t\leq 1$. Also, for $t<0$, there are singular inner functions that turn out to be shift-cyclic in $\mathcal{D}_{t}$. Korenblum [63] classified all shift-cyclic inner functions in $L^{2}_{a}(\mathbb{D})$. With these properties in mind, we highlight the following results from [24] (also see [13] for a modern treatment of Dirichlet-type spaces on $\mathbb{D}$).
@@
    If $f\in H^{2}(\mathbb{D})$ [^10] This result is valid for all $f\in H^{1}(\mathbb{D})$ as well, since one can check that $H^{1}(\mathbb{D})\subset L^{2}_{a}(\mathbb{D})$. then it is shift-cyclic in $L^{2}_{a}(\mathbb{D})$ if and only if its inner part is shift-cyclic in $L^{2}_{a}(\mathbb{D})$ (which are determined in [63]).
@@
Lastly, we remark that a polynomial is shift-cyclic in $\mathcal{D}_{t}$ with $t\leq 1$ if and only if it is non-vanishing on $\mathbb{D}$. This follows from the first part of (4) above, combined with [35, Theorem 9.2.1]. As noted before, a polynomial is shift-cyclic in $\mathcal{D}_{t}$ with $t>1$ if and only if it is non-vanishing on $\overline{\mathbb{D}}$. We saw with the Hardy spaces over $\mathbb{D}^{d}$ that determining the shift-cyclic polynomials is tractable for all $d\in\mathbb{N}$. In the following two subsections, we introduce generalizations of the Dirichlet-type spaces over $\mathbb{D}^{d}$ and $\mathbb{B}_{d}$, and note that the problem of determining shift-cyclic polynomials when $d=2$ is resolved, but it remains open for $d>2$.
@@ 3.2. On the unit polydisk $\mathbb{D}^{d}$
Moreover, we obtain the Bergman, Hardy and Dirichlet spaces on $\mathbb{D}^{d}$ by taking $t=-1,0$ and $1$ respectively. For $t>1$, $\mathcal{D}_{t}(\mathbb{D}^{d})$ is a subalgebra of functions in $\operatorname{Hol}(\mathbb{D}^{d})$ that are continuous up to $\overline{\mathbb{D}^{d}}$ and thus, is its own multiplier algebra. Also, just like in the one variable case, a function $f$ is shift-cyclic in $\mathcal{D}_{t}(\mathbb{D}^{d})$ with $t>1$ if and only if it is non-vanishing on $\overline{\mathbb{D}^{d}}$ (see [15, Section 1.1]).
While we can extend some of the general properties mentioned in the previous section to all $d>1$, one only needs to look at the polynomials to realize that classifying all shift-cyclic functions is further beyond the scope for these spaces than one might anticipate. In fact, we only have a complete classification of shift-cyclic polynomials for $d=2$. Let $\mathcal{Z}_{\Gamma}(P)$ denote the set of zeros of a polynomial $P\in\mathcal{P}_{d}$ on some $\Gamma\subseteq\mathbb{C}^{d}$. The following is a joint result of Bénéteau, Condori, Knese, Kosiński, Liaw, Seco and Sola [14, 15] [^11] See also [62] for a similar result for the anisotropic Dirichlet-type spaces on $\mathbb{D}^{2}$..
@@
If $P$ is an irreducible polynomial with $\mathcal{Z}_{\mathbb{D}^{2}}(P)=\emptyset$, then it is shift-cyclic in $\mathcal{D}_{t}(\mathbb{D}^{2})$ for
@@
    $t>1$ if and only if $\mathcal{Z}_{\mathbb{T}^{2}}(P)=\emptyset$.
There are a few things to unpack here. First, the proof of this theorem relies on something called a determinantal representation for polynomials that are non-vanishing on $\mathbb{D}^{2}$, which may not always exist for $d>2$ [60, 61]. Next, this result clearly indicates how boundary zeros are important to study shift-cyclicity of polynomials. In particular, the authors use the notion of Riesz $t$-capacity $\operatorname{cap}_{t}$ to determine the non-cyclicity of a polynomial $P$ in $\mathcal{D}_{t}(\mathbb{D}^{2})$ for large enough $t$, based on how large $\operatorname{cap}_{t}(\mathcal{Z}_{\mathbb{T}^{2}}(P))$ is.
Arguably, the most interesting aspect of Theorem 3.4 is the case $1/2<t\leq 1$, where finiteness of the boundary zero set is crucial to determining whether a polynomial is shift-cyclic or not. The authors achieve this via the use of tools from semi-analytic geometry. More specifically, they use an inequality of Łojasiewicz [67], which states that if $f$ is a non-zero real analytic function on some open set $U\subset\mathbb{R}^{N}$ with $N\in\mathbb{N}$ whose zero set $\mathcal{Z}_{U}(f)$ in $U$ is non-empty, then, for every compact $V\subset U$, there are constants $C_{V}>0$ and $n_{V}\in\mathbb{N}$ such that
@@
Applying this inequality to $|P|^{2}$, where $P$ is a non-vanishing polynomial on $\mathbb{D}^{2}$ with finitely many zeros in $\mathbb{T}^{2}$, allows us to compare $P$ with other polynomials that are shift-cyclic (see [15, Lemma 3.3]). This further enables us to determine the shift-cyclicity of such $P$ in $\mathcal{D}_{t}(\mathbb{D}^{2})$ when $t\leq 1$.
Bergqvist [16] extended some of these ideas to the $d>2$ case and presented several conditions on polynomials that guarantee shift-cyclicity as well as non-cyclicity of polynomials in $\mathcal{D}_{t}(\mathbb{D}^{d})$. However, the problem of determining shift-cyclicity of a general polynomial in $\mathcal{D}_{t}(\mathbb{D}^{d})$ remains unresolved.